OpenStudy (anonymous):

2√(x+2) -3=7 Solve for X , please sure all STEPS !!!!!!!

OpenStudy (anonymous):

Everything in (----) is under the square root

OpenStudy (anonymous):

SHOW ALL STEPS

OpenStudy (anonymous):

Generally, when solving for x what would you do first: addition/subtraction, or multiplication/division?

OpenStudy (opcode):

Solve for x over the real numbers: \(2 sqrt(x+2)-3 = 7\) Isolate terms with x to the left hand side. Add 3 to both sides: \(2 sqrt(x+2) = 10\) Divide both sides by a constant to simplify the equation. Divide both sides by 2: \(sqrt(x+2) = 5\) Eliminate the square root on the left hand side. Raise both sides to the power of two: \(x+2 = 25\) Solve for x. Subtract 2 from both sides: Answer: \( | x = 23 | \)

OpenStudy (anonymous):

-_- Fine. Or you could just read @Opcode 's solution and not think through it.

OpenStudy (opcode):

Sorry. He was spamming the chat -.-

OpenStudy (anonymous):

haha it's okay

OpenStudy (anonymous):

Oh my goodness, now I see what you mean.